# Dense Video Object Captioning from Disjoint Supervision

### v1

@@ B.3 CHOTA Evaluation code
We will release our CHOTA evaluation code. This evaluation code is based on the official HOTA implementation [^2] https://github.com/JonathonLuiten/TrackEval. The original code is under an MIT license.
B.4 Full training details
@@
As mentioned in Sec. 4.1.3, our model is based on GRiT wu2022grit . The original GRiT code [^3] https://github.com/JialianW/GRiT is released under an MIT license. Following GRiT, we use a ViTDet-Base dosovitskiy2020image ; li2022exploring backbone, a CenterNet zhou2019objects region proposal network and RoI Head, and a randomly-initialized text decoder following that of GIT wang2022git . The text decoder consists of 6 self-attention layers with casual feature masks wang2022git . All model architecture parameters follow the defaults from GRiT wu2022grit .
@@
The original GRiT uses an MAE pretrained checkpoint, while in our case we found a CLIP pretrained checkpoint radford2021learning performs better on our task. To fit more frames into memory for both training and evaluation, we use a $384\!\times\!384$ input size instead of the original $1024\!\times\!1024$. This choice moderately decreases dense image captioning performance on Visual Genome (from 17.3 $\text{AP}_{M}$to 15.7 $\text{AP}_{M}$).
@@
During disjoint multi-dataset pretraining, we sample batches from different datasets in an even ratio ($1:1:1:1$). For images datasets, a batch is composed of different images; for video datasets, we put the time dimension in batches and always guarantee images in the same mini-batch are from the same video. We use a local batch size of either 1 video (consisting of 8 sampled frames), or 8 images. As we use 32 TPU cores, this means that our global batch size is either 32 videos or 256 images. We use the AdamW optimizer with a learning rate of $2\times 10^{-4}$, weight decay of $0.05$, and a layerwise learning rate decay of $0.7$ li2022exploring ; wu2022grit . We train for $22.5\times 10^{3}$ iterations per dataset, decreasing the learning rate by a factor of 10 after $90\%$ and $97.5\%$ of the training schedule wu2022grit . For pretraining on all the 4 datasets in Sec. 3.4, this corresponds to a total of $90\times 10^{3}$ iterations, which took approximately 20 hours on 32, 16GB TPU cores.
@@
For VidSTG zhang2020does finetuning, we sample 16 frames in training, and run on all 200 frames in testing. For VLN voigtlaender2023connecting finetuning, we use the 3 annotated frames in both training and evaluation. For finetuning experiments on both datasets, we use a video batch size $16$ and train for $11.25\times 10^{3}$ iterations, with a learning rate of $10^{-5}$, weight decay of $0.05$, and layerwise-learning decay of $0.7$ li2022exploring . The finetuning took approximately 6 hours on 16, 16GB TPU cores for VidSTG, and about 2 hours on 16, 16GB TPU cores for VLN. Inference on VidSTG requires 32GB of TPU memory to fit 200 frames.
@@
The detection loss zhou2019objects involves a center heatmap loss, a bounding box regression loss, and a classification and bounding box refinement loss in the RoI head:
@@
$L_{heatmap}(Y,\bar{Y})=\frac{1}{n}\sum_{ij}\begin{cases}(1-{Y}_{ij})^{\alpha}\log({Y}_{ij})&\!\text{if}\ \bar{Y}_{ij}=1\\ (1-\bar{Y}_{ij})^{\beta}({Y}_{ij})^{\alpha}\log(1-{Y}_{ij})&\!\text{otherwise,}\end{cases}$ (3)
where $n$ is the number of objects in the image, $\alpha=2$ and $\beta=4$ are the focal loss weights lin2017focal .
$L_{reg}$ is a gIoU loss union2019metric :
$L_{reg}(B,\bar{B})=\frac{1}{n}\sum_{i}(\text{IoU}(B_{i},\bar{B}_{i})-\frac{|C_{i}\backslash(B_{i}\cup\bar{B}_{i})|}{|C_{i}|}),$ (4)
@@
$L_{assoc}=-\frac{1}{M}\sum_{ij}(\hat{A}_{ij}\log{A_{ij}}+(1-\hat{A}_{ij})\log{(1-A_{ij})})).$ (7)
The captioning loss is a softmax on each predicted word over the entire vocabulary, with a label smoothing co-efficient of $0.1$ following GIT wang2022git .
$L_{caption}=\frac{1}{L}\sum_{i=1}^{L}\text{Smooth-Softmax-CE}(\text{Decode}(f,\bar{y}_{1:i-1}),\bar{y}_{i}),$ (8)
@@
Appendix C Additional Experimental Analysis
C.1 Ablation of hard tracking sampling.
We analyze the effect of the number of sampled frames, $m$, in hard-aggregation (Sec. 3.3) in Tab. 7. With hard-aggregation, the captioning accuracy benefits from a larger number of frames $m$, thanks to longer input-sequence length. However this also costs more TPU memory in both training and testing. We use $m=6$ in our ablation experiments (Tab. 4) as it achieves a good balance between accuracy and computation. Moreover, it also follows the default number of frames used in the GIT wang2022git video captioning model. On the other hand, soft-aggregation (i.e., weighted-sum features) does not have this hyperparameter and has a constant memory cost, and achieves a higher caption accuracy. However, it also results in a higher number of caption switches.
Num sampled frames ($m$) CHOTA($\uparrow$) DetA($\uparrow$) AssA($\uparrow$) CapA($\uparrow$) $\text{AP}_{M}$($\uparrow$) CapSwitch($\downarrow$)
1 51.4 66.1 59.0 34.8 68.9 15.8
2 53.0 66.1 59.2 38.1 70.6 15.9
4 53.3 66.2 59.5 38.4 70.9 14.7
6 53.5 66.2 59.4 39.0 71.2 14.5
8 53.7 66.2 59.4 39.4 71.2 12.9
Soft aggregation 54.3 66.1 59.8 40.5 71.2 17.5
Table 7: Hyper-parameter sweep for number of sampled frames, $m$, for hard-tracking. We show results on VidSTG zhang2020does validation set. The models are based on #8 of Tab. 3 on VidSTG. Results with hard-feature aggregation get improved with more sampled frames, given the cost of increased training and testing memory.
C.2 Using the UVO dataset for disjoint pretraining
For the disjoint pretraining of our model (Sec. 3.4), we used Augmented COCO as our tracking dataset. Another alternative would have been to use UVO wang2021unidentified , which contains real-world videos, but is relatively small at only 5000 videos.
Table 8 compares Aug-COCO and UVO under the setting of Tab. 2 and Tab. 3 #8, both using a default multi-dataset sampling ratio $1:1:1:1$. We observe that disjoint pretraining with Aug-COCO consistently performs better than UVO in both zero-shot and finetuning scenarios, thus motivating our choice to use Aug-COCO for our experiments in the main paper.
@@ C.2 Using the UVO dataset for disjoint pretraining
Aug-COCO 54.3 66.1 59.8 40.5 71.2
Table 8: Results using UVO wang2021unidentified as the tracking dataset. We show both zero-shot results (top) and finetuning results (bottom) on VidSTG datasets. For reference, we also include our results of using Aug-COCO (#8 of Tab. 2 and Tab. 3). Aug-COCO performs better in both settings, motivating our choice to use it in the main paper.
C.3 Detailed captioning results
@@
4 10.2 7.3 8.7 14.7 9.1 3.7 9.1 14.3
5 9.4 7.3 8.1 12.9 7.7 3.0 7.7 12.5
6 4.0 1.9 6.1 4.0 6.4 2.5 8.2 8.4
7 9.8 7.4 8.4 13.7 8.2 3.0 8.3 13.3
8 10.0 7.1 8.5 14.4 9.4 4.3 9.4 14.5
Table 9: Detailed captioning metrics of our zero-shot evaluation (Tab. 2). We show the individual captioning metrics CIDEr, METEOR, and SPICE for each row on both datasets.
@@
0 35.8 43.2 29.0 35.0 8.8 4.2 14.8 7.4
1 36.8 46.1 28.5 35.8 8.2 6.7 13.9 3.9
2 38.0 47.5 29.8 36.7 16.6 13.6 20.1 16.1
3 33.0 35.8 28.7 35.8 14.5 9.4 19.6 14.5
4 39.1 49.3 30.3 37.6 17.6 13.9 21.2 17.7
5 39.6 50.2 30.5 38.1 17.2 14.8 20.4 16.4
6 37.1 46.5 28.9 35.7 11.6 8.1 17.2 9.5
7 40.7 52.9 30.8 38.4 17.6 14.2 21.5 17.1
8 40.5 52.6 30.4 38.5 17.6 14.1 21.4 17.4
Table 10: Detailed captioning metrics of our finetuning evaluation (Tab. 3). We show the individual captioning metrics CIDEr, METEOR, and SPICE for each row on both datasets.
@@
STGRN zhang2020does - - - 38.0 48.5 19.8
STVGBert su2021stvgbert - - - 47.3 - 24.0
TubeDETR yang2022tubedetr 56.4 47.2 28.7 59.0 48.1 30.4
STCAT jin2022embracing - - - 61.7 50.8 33.1
@@
C.4 VidSTG spatio-temporal grounding evaluation
Table 6 of the main paper comapred to prior methods on the spatial-video grounding task on VidSTG (where the input videos were assumed to already be temporally trimmed). In Tab. 11, we report results for spatial-, temporal- and spatio-temporal grounding by reporting the Spatial IoU (sIoU), Temporal IoU (tIoU) and Video IoU (vIoU) respectively.
The sIoU assumes the video is already temporally trimmed before evaluation, thus evaluating spatial localization. Similarly, the tIoU assumes that the video is already cropped spatially around the object of interest, and only the temporal extent of the query sentence needs to be determined, thereby evaluating temporal localization. The vIoU that evaluates both spatial and temporal localization.
Our model was designed for the Dense VOC task, and not grounding, and we were able to perform grounding by selecting the bounding boxes with the highest likelihood of generating the target sentence (Sec. 3.5). As shown in Tab. 11, this approach works well for spatial-grounding, outperforming prior works in terms of the sIoU. However, as our model first generates object trajectories, without taking the input query into account, it struggles more at temporal localization. Nevertheless, it still achieves competitive results compared to prior works for both the tIoU and vIoU, although our model was not designed specifically for this task like the other methods in Tab. 11 which include explicit temporal localization modules within the network architecture.

### v2

@@ 0.A.2 CHOTA Details and Evaluation code
We attach our CHOTA evaluation code as chota.py. This evaluation code is based on the official HOTA implementation***https://github.com/JonathonLuiten/TrackEval. The original code is under an MIT license.
0.A.3 Full training details
@@
As mentioned in Sec. References, our model is based on GRiT [65]. The original GRiT code***https://github.com/JialianW/GRiT is released under an MIT license. Following GRiT, we use a ViTDet-Base [22, 39] backbone, a CenterNet [88] region proposal network and RoI Head, and a randomly-initialized text decoder following that of GIT [61]. The text decoder consists of 6 self-attention layers with casual feature masks [61]. All model architecture parameters follow the defaults from GRiT [65].
@@
The original GRiT uses an MAE pretrained checkpoint, while in our case we found a CLIP pretrained checkpoint [48] performs better on our task. To fit more frames into memory for both training and evaluation, we use a $384\!\times\!384$ input size instead of the original $1024\!\times\!1024$. This choice moderately decreases dense image captioning performance on Visual Genome (from 17.3 $\text{AP}_{M}$to 15.7 $\text{AP}_{M}$).
@@
During disjoint multi-dataset pretraining, we sample batches from different datasets in an even ratio ($1:1:1:1$). For image datasets, a batch is composed of different images; for video datasets, we put the time dimension in batches and always guarantee images in the same mini-batch are from the same video. We use a local batch size of either 1 video (consisting of 8 sampled frames), or 8 images. As we use 32 TPUs, this means that our global batch size is either 32 videos or 256 images. We use the AdamW optimizer with a learning rate of $2\times 10^{-4}$, weight decay of $0.05$, and a layerwise learning rate decay of $0.7$ [39, 65]. We train for $22.5\times 10^{3}$ iterations per dataset, decreasing the learning rate by a factor of 10 after $90\%$ and $97.5\%$ of the training schedule [65]. For pretraining on all the 4 datasets in Sec. References, this corresponds to a total of $90\times 10^{3}$ iterations, which took approximately 20 hours on 32, 16GB TPUs cores.
@@
For VidSTG [83] finetuning, we sample 16 frames in training, and run on all 200 frames in testing. For VLN [60] finetuning, we use the 3 annotated frames in both training and evaluation. For finetuning experiments on both datasets, we use a video batch size $16$ and train for $11.25\times 10^{3}$ iterations, with a learning rate of $10^{-5}$, weight decay of $0.05$, and layerwise-learning decay of $0.7$ [39]. The finetuning took approximately 6 hours on 16, 16GB TPUs for VidSTG, and about 2 hours on 16, 16GB TPUs for VLN. Inference on VidSTG requires 32GB of TPU memory to fit 200 frames.
@@
The detection loss [88] involves a center heatmap loss, a bounding box regression loss, and a classification and bounding box refinement loss in the RoI head:
@@
$\scriptsize L_{heatmap}(Y,\bar{Y})=\frac{1}{n}\sum_{ij}\begin{cases}(1-{Y}_{ij})^{\alpha}\log({Y}_{ij})&\!\text{if}\ \bar{Y}_{ij}=1\\ (1-\bar{Y}_{ij})^{\beta}({Y}_{ij})^{\alpha}\log(1-{Y}_{ij})&\!\text{otherwise,}\end{cases}$ (5)
where $n$ is the number of objects in the image, $\alpha=2$ and $\beta=4$ are the focal loss weights [40].
$L_{reg}$ is a gIoU loss [58]:
$\footnotesize L_{reg}(B,\bar{B})=\frac{1}{n}\sum_{i}(\text{IoU}(B_{i},\bar{B}_{i})-\frac{|C_{i}\backslash(B_{i}\cup\bar{B}_{i})|}{|C_{i}|}),$ (6)
@@
$\footnotesize L_{assoc}=-\frac{1}{M}\sum_{ij}(\hat{A}_{ij}\log{A_{ij}}+(1-\hat{A}_{ij})\log{(1-A_{ij})})).$ (9)
The captioning loss is a softmax on each predicted word over the entire vocabulary, with a label smoothing co-efficient of $0.1$ following GIT [61].
$L_{caption}=\frac{1}{L}\sum_{i=1}^{L}\text{CE}(\text{Decode}(f,\bar{y}_{1:i-1}),\bar{y}_{i}),$ (10)
@@
Appendix 0.B Additional Experimental Analysis
0.B.1 Ablation of hard tracking sampling.
We analyze the effect of the number of sampled frames, $m$, in hard-aggregation (Sec. References) in Tab. 7. With hard-aggregation, the captioning accuracy benefits from a larger number of frames $m$, thanks to longer input-sequence length. However, this also costs more TPU memory in both training and testing. We use $m=6$ in our ablation experiments (Tab. References) as it achieves the best accuracy. It also follows the default number of frames used in the GIT [61] video captioning model.
$m$ CHOTA($\uparrow$) DetA($\uparrow$) AssA($\uparrow$) CapA($\uparrow$) $\text{AP}_{M}$($\uparrow$)
1 53.6 64.3 66.3 36.1 68.7
2 54.1 64.3 66.2 37.3 68.7
4 54.4 64.3 65.7 37.8 68.7
6 54.9 64.2 65.9 39.1 69.1
8 54.6 64.3 66.1 38.4 69.1
Table 7: Hyper-parameter sweep for number of sampled frames, $m$, for hard-tracking. We show results on VidSTG [83] validation set. The models are based on #2 of Tab. References on VidSTG. Results with hard-feature aggregation get improved with more sampled frames and get saturated with 6 frames.
@@ Appendix 0.B Additional Experimental Analysis
Aug-COCO 56.9 65.8 70.4 39.7 71.5
Table 8: Results using UVO [63] as the tracking dataset. We show both zero-shot results (top) and finetuning results (bottom) on VidSTG datasets. For reference, we also include our results of using Aug-COCO (#8 of Tab. References and Tab. References). Aug-COCO performs better in both settings, motivating our choice in the main paper.
0.B.2 Using the UVO dataset for disjoint pretraining
For the disjoint pretraining of our model (Sec. References), we used Augmented COCO as our tracking dataset. Another alternative would have been to use UVO [63], which contains real-world videos, but is relatively small at only 5000 videos.
Table 8 compares Aug-COCO and UVO under the setting of Tab. References and Tab. References #8, both using a default multi-dataset sampling ratio $1:1:1:1$. We observe that disjoint pretraining with Aug-COCO consistently performs better than UVO in both zero-shot and finetuning scenarios, thus motivating our choice to use Aug-COCO for our experiments in the main paper.
0.B.3 Detailed captioning results
@@
4 8.5 4.4 7.7 13.4 8.5 3.1 8.7 13.8
5 8.1 4.0 7.2 13.0 7.8 3.1 7.8 12.4
6 4.9 3.3 7.2 4.2 7.5 3.7 9.4 9.6
7 9.1 5.2 8.3 13.7 9.0 3.9 9.2 13.9
8 9.8 7.0 9.1 13.3 9.7 4.6 9.9 14.6
Table 9: Detailed captioning metrics of our zero-shot evaluation (Tab. References). We show the individual captioning metrics CIDEr, METEOR, and SPICE for each row on both datasets.
@@
0 35.8 43.2 29.0 35.0 8.7 3.9 14.8 7.3
1 34.8 41.7 27.5 35.5 8.2 6.7 14.0 3.9
2 39.1 49.4 30.3 37.6 16.7 13.9 20.1 16.1
3 31.6 33.8 27.2 33.6 14.5 9.6 19.6 14.2
4 39.2 49.9 30.3 37.5 17.8 13.9 21.4 17.9
5 38.4 48.3 30.0 36.9 17.4 15.1 20.5 16.5
6 38.8 49.6 29.9 37.1 11.6 8.2 17.3 9.5
7 40.1 51.5 30.8 38.1 17.7 14.6 21.5 17.1
8 39.7 51.0 30.6 37.7 17.7 14.3 21.5 17.5
Table 10: Detailed captioning metrics of our finetuning evaluation (Tab. References). We show the individual captioning metrics CIDEr, METEOR, and SPICE for each row on both datasets.
@@
STGRN [83] - - - 38.0 48.5 19.8
STVGBert [56] - - - 47.3 - 24.0
TubeDETR [72] 56.4 47.2 28.7 59.0 48.1 30.4
STCAT [31] - - - 61.7 50.8 33.1
@@
0.B.4 VidSTG spatio-temporal grounding evaluation
Table References of the main paper compared to prior methods on the spatial-video grounding task on VidSTG (where the input videos were assumed to already be temporally trimmed). In Tab. 11, we report results for spatial-, temporal- and spatio-temporal grounding by reporting the Spatial IoU (sIoU), Temporal IoU (tIoU) and Video IoU (vIoU) respectively.
The sIoU assumes the video is already temporally trimmed before evaluation, thus evaluating spatial localization. Similarly, the tIoU assumes that the video is already cropped spatially around the object of interest, and only the temporal extent of the query sentence needs to be determined, thereby evaluating temporal localization. The vIoU evaluates both spatial and temporal localization.
Our model was designed for the Dense VOC task, and not grounding, and we were able to perform grounding by selecting the bounding boxes with the highest likelihood of generating the target sentence (Sec. References). As shown in Tab. 11, this approach works well for spatial-grounding, outperforming prior works in terms of the sIoU. However, as our model first generates object trajectories, without taking the input query into account, it struggles more at temporal localization. Nevertheless, it still achieves competitive results compared to prior works for both the tIoU and vIoU, although our model was not designed specifically for this task like the other methods in Tab. 11 which include explicit temporal localization modules within the network.
Appendix 0.C Broader impact and potential negative impact
Our work presents a new task and model for dense video object captioning. This task represents a general technology with a wide range of potential applications. Whilst we are unaware of all potential applications of such models, it is important to be cognizant that each application has its own merits and societal implications depending on the intentions of the individuals building and using the system. For example, we believe that the Dense VOC models can be used as part of systems to improve video search and retrieval, though they could also be used in video surveillance systems too. Additionally, we note that training datasets, especially for captioning [28, 84], can contain biases that may render models trained on them unsuitable for deployment.